A source of sound S having frequency f. Wind is blowing from source to observer O with velocity u. If speed of sound with respect to air is C, the wavelength of sound detected by O is:

Grade:12

3 Answers

Arun
6 years ago
Dear Raj
I am getting my answer as
(C + u) /f
Please let me know, if this correct then I will share my solution.
Regards
Arun (askIITians forum expert)
Bhavyajain
6 years ago
Yes your answer is correctBut how do you do this question?Please share your solution as soon as possible
Lovish Vijayvargiya
5 years ago
Velocity of source from observer = V-u-C
V= velocity of sound
F= (v-0/v-(v-u-C)) * f
F= (v/c+u)*f
v=(v/c+u)*f*lambda
Lambda= c+u/f